10. Consider the forced but undamped system described by the initial value prob- lem u + 25u = sin(wt), (0) = 0, ta (0) =1.
Answer #1

U²+25 u = sinot Le col=0 L(0,) a the auxilary Equation.be m m+25=0 m=tsi General > Solun Ug = G Gast +Csinst Particular Inti1+ U() = (es way Requared solun (25-ww] Sinst thinieot 5 b 7 when w-4, 4.5, 4.9. o u(t) 25-16-4 1 sinst thin4. sinst tllingt t2when le (+) W = =5 undefined & Snii 25-co a But limeting Graph not exist ult) : is - w-w) finst + sinof? (25-60) here Wo is b3III 0:5 -0.5 0.5 -0.5 + 1 Х sin(52)+* sin(4x) 9 21 no graph4:20 PM ...1.0KB/S @MO 33 -1 0 -1- + 1 Х 2 Х -391 495 sin(52)+ 99 100 sin(4 3 powered by desmos2no graph= 1 0 -1 + 1 Х 2 Х -391 495 sin(5x)+ 99 100 sin(4 3.3no graph